OpenStudy (anonymous):

Determine dy/dx for the following: x cos(2x^2)

OpenStudy (anonymous):

use product rule and along the way the chain rule. \[\frac{d}{dx}\cos(2x^2)=\cos(2x^2)-4x\sin(2x^2)\]